# LITP-25-17

### v1

@@ 8 Summary and Conclusions
In conclusion, long–range forces can be incorporated into the EFT–hedron without disrupting the UV–IR link between low–energy EFTs and their microscopic completions. An especially promising direction is the extension of the SS–matrix bootstrap [Paulos:2016but, Paulos:2017fhb, Cordova:2019lot, EliasMiro:2019kyf, Homrich:2019cbt, Karateev:2019ymz, Guerrieri:2020bto, EliasMiro:2021nul, Guerrieri:2021ivu, Kruczenski:2022lot, EliasMiro:2022xaa, He:2023lyy, Guerrieri:2023qbg, Guerrieri:2024ckc, Eckner:2024ggx, Eckner:2024pqt, Gumus:2024lmj, He:2025gws], which in $D=4$ currently operates only under the extreme assumption that both $\alpha$ ($G$) and $\alpha_{\mathcal{E}}$ ($G_{\mathcal{E}}$) vanish, or the SS–matrix becomes trivial. With stripped amplitudes, we can now move away from this corner and allow for finite $\alpha_{\mathcal{E}}$ ($G_{\mathcal{E}}$), thereby incorporating the leading long–range effects directly into numerical bootstrap analyses. This is particularly important in the case of electromagnetism for realistic applications of positivity and bootstrap constraints to particle phenomenology. Moreover, as already remarked, the future challenge is to develop a approach enabling to include corrections beyond the leading-log approximation, possibly based on IR–finite and physically relevant observables.
@@ Appendix A Symmetric Dispersions
We give here further details on the extension of tutu–symmetric dispersion relations [Bellazzini:2025shd] to the case of massive particles. These dispersion relations are used in the main text to bound $\pi^{+}\pi^{+}$ scattering in Section 6.2.
@@
In this section we provide a brief review of the stustu–symmetric dispersion relations for massless scattering [Li:2023qzs], which are used in Section 7 to obtain bounds on gravitational scattering. It is convenient to work in the variables
@@
where the contour is represented for instance on Fig. 6 of [Beadle:2025cdx].
@@ B.1 $\pi^{+}\pi^{0}$ Scattering
This removes any contribution from Wilson coefficients with powers of $q$ larger than two, i.e. any coeffients in the arc other than $c_{2,0}$ and $c_{3,1}$. We then use semidefinite optimization [Simmons-Duffin:2015qma] to optimize the $a_{i}$ coefficients to get bounds on the remaining Wilson coefficients ($c_{2,0}$ and $c_{3,1}$ in the example above). In order to have enough free parameters to remove contributions from the higher Wilson coefficients, get the optimal bound on the ones we do not kill, and require positivity in the UV, we observe that $k_{\text{max}}=N+2$. Finally, we evaluate the bounds as we increase $N$ (and $k_{\text{max}}$), and check if the bounds converge. For the tree-level amplitude in (6.4), the convergence of the bound $c_{3,1}/c_{2,0}$ as we increase $N$ is given in Fig. 7.
@@ B.2 $\pi^{+}\pi^{+}$ and $\pi^{0}\pi^{0}$ Scattering
The process is exactly the same for gravity, where we can again analytically evaluate the necessary integrals in the scaling limit. We use a slightly different set of basis functions which matches the choice of [Caron-Huot:2021rmr, Chang:2025cxc, Beadle:2025cdx].
@@ C.1 One-loop stripped amplitudes
$\begin{split}\mathcal{M}^{\text{1-loop}}_{\mathcal{E}}(s,t){\Big|}^{\mathrm{GR}}&=-\frac{G\hat{c}_{2,0}}{\pi}\Bigg[\frac{1}{150}\left(75\pi^{2}-6346\right)stu-\frac{1}{900}\left(150\pi^{2}-4697\right)\left(s^{3}+t^{3}+u^{3}\right)\\ &+\left(s^{3}-stu\right)\log^{2}\left(-\frac{s}{\hat{\mathcal{E}}^{2}}\right)+\left(t^{3}-stu\right)\log^{2}\left(-\frac{t}{\hat{\mathcal{E}}^{2}}\right)+\left(u^{3}-stu\right)\log^{2}\left(-\frac{u}{\hat{\mathcal{E}}^{2}}\right)\\ &+\frac{1}{30}\log\left(-\frac{s}{\hat{\mu}_{\text{UV}}^{2}}\right)\left(181stu-76s^{3}\right)+\frac{1}{30}\left(181stu-76t^{3}\right)\log\left(-\frac{t}{\hat{\mu}_{\text{UV}}^{2}}\right)\\ &+\frac{1}{30}\left(181stu-76u^{3}\right)\log\left(-\frac{u}{\hat{\mu}_{\text{UV}}^{2}}\right)\Bigg]\,.\end{split}$ (C.4)
@@ C.2 Various Arcs
$\begin{split}\mathcal{C}_{0}(q)&=\frac{\hat{c}_{2,0}G}{\pi}\Bigg[-\frac{q^{2}}{30M^{2}}\left(181+60\log(\mathcal{E}^{2}/M^{2})\right)\log\left(\frac{M^{2}}{q^{2}}+1\right)\\ &-4\log(\mathcal{E}^{2}/M^{2})-\frac{533q^{2}}{20M^{2}}-\frac{136}{15}+2\frac{q^{2}}{M^{2}}\,\text{Li}_{2}\left(-\frac{M^{2}}{q^{2}}\right)-\frac{21q^{2}}{2M^{2}}\log(\mu_{\text{UV}}/M^{2})\Bigg]\,.\end{split}$ (C.11)

### v2

@@ 8 Summary and Conclusions
In conclusion, long–range forces can be incorporated into the EFT–hedron without disrupting the UV–IR link between low–energy EFTs and their microscopic completions. An especially promising direction is the extension of the SS–matrix bootstrap [123, 124, 57, 76, 101, 106, 86, 77, 88, 108, 74, 96, 85, 84, 73, 72, 89, 97], which in $D=4$ currently operates only under the extreme assumption that both $\alpha$ ($G$) and $\alpha_{\mathcal{E}}$ ($G_{\mathcal{E}}$) vanish, or the SS–matrix becomes trivial. With stripped amplitudes, we can now move away from this corner and allow for finite $\alpha_{\mathcal{E}}$ ($G_{\mathcal{E}}$), thereby incorporating the leading long–range effects directly into numerical bootstrap analyses. This is particularly important in the case of electromagnetism for realistic applications of positivity and bootstrap constraints to particle phenomenology. Moreover, as already remarked, the future challenge is to develop a approach enabling to include corrections beyond the leading-log approximation, possibly based on IR–finite and physically relevant observables.
@@ Appendix A Symmetric Dispersions
We give here further details on the extension of tutu–symmetric dispersion relations [25] to the case of massive particles. These dispersion relations are used in the main text to bound $\pi^{+}\pi^{+}$ scattering in Section 6.2.
@@
In this section we provide a brief review of the stustu–symmetric dispersion relations for massless scattering [111], which are used in Section 7 to obtain bounds on gravitational scattering. It is convenient to work in the variables
@@
where the contour is represented for instance on Fig. 6 of [17].
@@ B.1 $\pi^{+}\pi^{0}$ Scattering
This removes any contribution from Wilson coefficients with powers of $q$ larger than two, i.e. any coeffients in the arc other than $c_{2,0}$ and $c_{3,1}$. We then use semidefinite optimization [132] to optimize the $a_{i}$ coefficients to get bounds on the remaining Wilson coefficients ($c_{2,0}$ and $c_{3,1}$ in the example above). In order to have enough free parameters to remove contributions from the higher Wilson coefficients, get the optimal bound on the ones we do not kill, and require positivity in the UV, we observe that $k_{\text{max}}=N+2$. Finally, we evaluate the bounds as we increase $N$ (and $k_{\text{max}}$), and check if the bounds converge. For the tree-level amplitude in (6.4), the convergence of the bound $c_{3,1}/c_{2,0}$ as we increase $N$ is given in Fig. 7.
@@ B.2 $\pi^{+}\pi^{+}$ and $\pi^{0}\pi^{0}$ Scattering
The process is exactly the same for gravity, where we can again analytically evaluate the necessary integrals in the scaling limit. We use a slightly different set of basis functions which matches the choice of [49, 53, 17].
@@ C.1 One-loop stripped amplitudes
$\begin{split}\mathcal{M}^{\text{1-loop}}_{\mathcal{E}}(s,t){\Big|}^{\mathrm{GR}}&=-\frac{G\hat{c}_{2,0}}{2\pi}\Bigg[\frac{1}{150}\left(75\pi^{2}-6346\right)stu-\frac{1}{900}\left(150\pi^{2}-4697\right)\left(s^{3}+t^{3}+u^{3}\right)\\ &+\left(s^{3}-stu\right)\log^{2}\left(-\frac{s}{\hat{\mathcal{E}}^{2}}\right)+\left(t^{3}-stu\right)\log^{2}\left(-\frac{t}{\hat{\mathcal{E}}^{2}}\right)+\left(u^{3}-stu\right)\log^{2}\left(-\frac{u}{\hat{\mathcal{E}}^{2}}\right)\\ &+\frac{1}{30}\log\left(-\frac{s}{\hat{\mu}_{\text{UV}}^{2}}\right)\left(181stu-76s^{3}\right)+\frac{1}{30}\left(181stu-76t^{3}\right)\log\left(-\frac{t}{\hat{\mu}_{\text{UV}}^{2}}\right)\\ &+\frac{1}{30}\left(181stu-76u^{3}\right)\log\left(-\frac{u}{\hat{\mu}_{\text{UV}}^{2}}\right)\Bigg]\,.\end{split}$ (C.4)
@@ C.2 Various Arcs
$\begin{split}\mathcal{C}_{0}(q)&=\frac{\hat{c}_{2,0}G}{2\pi}\Bigg[-\frac{q^{2}}{30M^{2}}\left(181+60\log(\mathcal{E}^{2}/M^{2})\right)\log\left(\frac{M^{2}}{q^{2}}+1\right)\\ &-4\log(\mathcal{E}^{2}/M^{2})-\frac{533q^{2}}{20M^{2}}-\frac{136}{15}+2\frac{q^{2}}{M^{2}}\,\text{Li}_{2}\left(-\frac{M^{2}}{q^{2}}\right)-\frac{21q^{2}}{2M^{2}}\log(\mu_{\text{UV}}/M^{2})\Bigg]\,.\end{split}$ (C.11)